NURBS today G. Farin

Programming assignment #4

due 4-30

Compare curvature plots of degree 2,3,4 B-spline curves. Pick your control points on parts of a spiral as one example. Since you will need first and second derivatives, you might want to implement a sufficiently general blossom routine.